OpenStudy (anonymous):

what is the binomial difference of (7n+(-2))-(8n+4)

OpenStudy (anonymous):

first thing you do is distribute negative sign. then group the same terms and calculate once you distribute negative sign, it's gonna be 7n-2-8n-4 keep in mind that you will be calculating based on the same expressions/terms hence, you will be calculating n^1 terms and n^0 terms so the answer will have two terms

OpenStudy (anonymous):

can u open iht

OpenStudy (anonymous):

7n - 2 - 8n - 4 = 7n - 8n - 2 - 4 = -n -6

OpenStudy (anonymous):

can't seem to open that file

OpenStudy (anonymous):

what kind of file is it? never seen .mht extension before

OpenStudy (anonymous):

i hv opened it.....

OpenStudy (anonymous):

thanks harkirat. & msayer3 thanks also!

OpenStudy (anonymous):

what should i do with the file.... it has opened up a test site of Plato Learning....

OpenStudy (anonymous):

I always get here too late

OpenStudy (anonymous):

jacyy....took the test got 9 out of ten correct.....☺

OpenStudy (anonymous):

I got them all right :) 10/10

OpenStudy (anonymous):

That was fun jacy, post more of those...